OpenStudy (anonymous):

f(t)=1.8(1.452)^t t=0 correspond to year2004. In what year will sales reach 20 billion?

OpenStudy (anonymous):

you are trying to solve \[20=1.8(1.452)^t\] for t. divide by 1.8 \[\frac{20}{1.8}=(1.452)^t\] take the log of both sides \[\ln(\frac{20}{1.8})=\ln((1.452)^t)=t\ln(1.452)\]

OpenStudy (anonymous):

solve for t: \[t=\frac{\ln(\frac{20}{1.8})}{\ln(1.452)}\]

OpenStudy (anonymous):

thats what i got, but the answer would be 6.5 so the year 2010,

OpenStudy (anonymous):

i get 6.4 so 6.4 years after 2004 is 2010.4 of there abouts.

OpenStudy (anonymous):

but the question before asked what are sales in 2010 and my answer was 16.9.....not 20. im confused

OpenStudy (anonymous):

maybe be safe and say by 2011

OpenStudy (anonymous):

i guess the additional .5 counts! let me try it.

OpenStudy (anonymous):

thanks so much

OpenStudy (anonymous):

welcome!